OpenStudy (anonymous):

what is the answer to -3√5+2√5

OpenStudy (jdoe0001):

\(\bf -3\sqrt{5}+2\sqrt{5}\qquad let's \qquad \sqrt{5}={\color{red}{ a}}\qquad thus\\ \quad \\ -3{\color{red}{ a}}+2{\color{red}{ a}}\)

OpenStudy (jdoe0001):

so, subtract that, see what you get

OpenStudy (the_fizicx99):

They have the same radicands so just -3 + 2 -3 + 2 = -1 \[-1\sqrt{5}\]

OpenStudy (anonymous):

-1√5 is the answer.